# Automated Privacy Information Annotation in Large Language Model Interactions

### v1

@@ Appendix I Discussion of Main Results
[privacy category 1, privacy category 2, … ]
Prompts for categories deduplication (Chinese).
[…]
{CJK}UTF8gbsn 给定以下<BLOCK-SIZE>个隐私类别及每个隐私类别对应的示例短语列表:
{CJK}UTF8gbsn <|INPUT|>
{CJK}UTF8gbsn
{CJK}UTF8gbsn 对隐私类别进行去重,如果隐私类别A和隐私类别B具有相同的含义,只保留类别A
{CJK}UTF8gbsn 让我们一步一步地思考,最后输出一个json列表。
{CJK}UTF8gbsn 以列表格式输出去重后的新的隐私类别,例如:
{CJK}UTF8gbsn [隐私类别1, 隐私类别2, … ]
Table 13: Prompt template for privacy phrase extraction with categories in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|CATEGORIES|> is the input of privacy categories.
Prompts for privacy phrase extraction (English).
[…]
A user poses a query
"<|INPUT|>"
to you.
@@
Give your think steps. The reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
@@
Privacy Categories: <|CATEGORIES|>
@@
    "phrase 1" : "privacy category 3",
    "phrase 2" : "privacy category 4"
Here are some examples
User’s Query: "I’m a teacher at China, majoring in computer science. I will go back home to London."
Privacy Categories: ["Name", "Age", "Gender", "Job", "Subject", "Location"]
@@
“‘json
    "teacher": "Job",
    "computer science": "Subject",
    "China": "Location",
    "London": "Location"
…… (other examples)
Prompts for privacy phrase extraction (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn 从请求的内容中提取短语,该短语明确揭示用户的,或与用户相关的人或事物的隐私敏感信息。
{CJK}UTF8gbsn 短语必须与请求内容中出现的完全一致。不要更改短语的格式或措辞,即拼写错误、空格数量、大小写等,都必须与请求内容完全相同。
{CJK}UTF8gbsn 给出你的思考步骤(Think Step by Step)。推理必须依赖于请求中包含的直接证据,不能基于假设
{CJK}UTF8gbsn 然后,将每个提取的短语与以下预定义列表中最合适的隐私类别匹配,并且如果提取的短语未与任何给定的隐私类别匹配,则应将其排除。
{CJK}UTF8gbsn
{CJK}UTF8gbsn 隐私类别:<|CATEGORIES|>
{CJK}UTF8gbsn
{CJK}UTF8gbsn 结果输出应为json格式,格式如下:
[…]
{CJK}UTF8gbsn     "短语1" : "隐私类别3",
{CJK}UTF8gbsn     "短语2" : "隐私类别4"
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一些示例
{CJK}UTF8gbsn 用户的请求:"我是中国的一名教师,主修计算机科学。我将回家到伦敦。"
{CJK}UTF8gbsn 隐私类别:["姓名", "年龄", "性别", "职业", "专业", "地点"]
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "教师": "职业",
{CJK}UTF8gbsn     "计算机科学": "专业",
{CJK}UTF8gbsn     "中国": "地点",
{CJK}UTF8gbsn     "伦敦": "地点"
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn …… (other examples)
Table 14: Prompt template for privacy phrase deduplication in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASES|> is the input of extracted privacy phrase.
Prompts for privacy phrase deduplication (English).
[…]
Given a user’s query and privacy phrases extracted from the query,
Query: "<|INPUT|>"
Privacy Phrases: <|PHRASES|>
@@
[Rule 2] If privacy phrase A is part of privacy phrase B, keep phrase which has more information and is clear.
@@
“‘json
[new privacy phrase 1, new privacy phrase 2, ]
Here is a examples:
Query: "We are happy and plan to play football."
Privacy Phrases: ["happy", "We are happy", "plan to play football", "football", "play football"]
@@
“‘json
["happy", "play football"]
Prompts for privacy phrase deduplication (Chinese).
[…]
{CJK}UTF8gbsn 给定一个用户请求和从请求中提取的隐私短语,
{CJK}UTF8gbsn 用户请求:"<|INPUT|> "
{CJK}UTF8gbsn 隐私短语:<|PHRASES|>
{CJK}UTF8gbsn
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step),严格按照给定的规则对隐私短语进行去重
{CJK}UTF8gbsn [规则1] 如果隐私短语A和隐私短语B具有相同的含义,保留更简洁明了的短语。
{CJK}UTF8gbsn [规则2] 如果隐私短语A是隐私短语B的一部分,保留信息量更大且清晰的短语。
{CJK}UTF8gbsn 给出你的思考步骤。然后以json列表格式输出去重后的新隐私短语,例如:
{CJK}UTF8gbsn “‘json
{CJK}UTF8gbsn [新隐私短语1, 新隐私短语2, ]
Table 15: Prompt template for filtering privacy phrase with rule 1 in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASE|> is the input of the phrase.
Prompts for filtering privacy phrase with rule 1 (English).
[…]
A user poses a query
"<|INPUT|>"
@@
Give your think steps. The reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
@@
“‘json
    "reason" : reason for your judgment,
    "judgment" : true or false
Here are some examples:
User’s Query: "We plan to play football with Ross."
Phrase: "play football with Ross"
@@
“‘json
    "reason" : "Playing football is the user’s plan that related to the user. Ross is also related to the user.",
    "judgment" : true
…… (other examples)
Prompts for filtering privacy phrase with rule 1 (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn 并给定用户的真实的身份标识
{CJK}UTF8gbsn 给定从请求中提取出短语如下:
{CJK}UTF8gbsn 短语:" <|PHRASE|> "
{CJK}UTF8gbsn
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step),判断该短语是否满足以下规则
{CJK}UTF8gbsn [规则] 该短语在用户请求的上下文中直接与用户或与用户相关的人或事物相关联。
{CJK}UTF8gbsn 给出你的思考步骤。推理必须依赖于请求中包含的直接证据,不能基于假设
{CJK}UTF8gbsn 然后,输出如下格式的json:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason" : 你判断的理由,
{CJK}UTF8gbsn     "judgment" : true 或 false
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:"我们计划和罗斯一起踢足球。"
{CJK}UTF8gbsn 短语:"和罗斯一起踢足球"
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason" : "踢足球是用户的计划,与用户相关。罗斯也与用户相关。",
{CJK}UTF8gbsn     "judgment" : true
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn …… (other examples)
Table 16: Prompt template for filtering privacy phrase with rule 2 in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASE|> is the input of the phrase.
Prompts for filtering privacy phrase with rule 2 (English).
[…]
A user poses a query
"<|INPUT|>"
@@
[Rule] the phrase should not be a general phrase that provide NO information related to the user based on the context of the user’s query.
Give your think steps. The reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
@@
“‘json
    "reason" : reason for your judgment,
    "judgment" : true or false
Here are some examples:
User’s Query: "I want to go a place."
Phrases: "place"
@@
“‘json
    "reason" : "The ‘place’ term has no clear reference.",
    "judgment" : false
…… (other examples)
Prompts for filtering privacy phrase with rule 2 (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
"<|INPUT|>"
{CJK}UTF8gbsn 并给定用户的真实的身份标识
{CJK}UTF8gbsn 给定从请求中提取出短语如下:
{CJK}UTF8gbsn 短语:" <|PHRASE|> "
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step),判断该短语是否满足以下规则
{CJK}UTF8gbsn [规则] 该短语不应是一个在用户请求上下文中不提供与用户相关信息的一般短语。
{CJK}UTF8gbsn 给出你的思考步骤。推理必须依赖于查询中包含的直接证据,不能基于假设
{CJK}UTF8gbsn 然后,输出如下格式的json:
“‘json
    "reason" : {CJK}UTF8gbsn 你判断的理由 ,
    "judgment" : true {CJK}UTF8gbsn 或 false
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:"我想去一个地方。"
{CJK}UTF8gbsn 短语:"地方"
{CJK}UTF8gbsn JSON输出:
“‘json
{CJK}UTF8gbsn     "reason" : "‘地方’一词没有明确的指代。" ,
    "judgment" : false
…… (other examples)
Table 17: Prompt template for privacy information annotation in step 4 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASE|> is the input of the phrase.
Prompts for privacy information annotation (English).
[…]
A user poses a query
"<|INPUT|>"
@@
    "phrase": PHRASE,
    "privacy information": PRIVACY-INFORMATION
Here are some examples:
User’s Query: "I am preparing to go back home to Minhang District, Shanghai."
Phrase: "Minhang District, Shanghai"
@@
“‘json
    "phrase": "No. 800, Dongchuan Road, Minhang District, Shanghai",
    "privacy information": "The user’s home address is No. 800, Dongchuan Road, Minhang District, Shanghai."
…… (other examples)
Prompts for privacy information annotation (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
"<|INPUT|>"
{CJK}UTF8gbsn 并给定用户的真实的身份标识
{CJK}UTF8gbsn 给定从请求中提取出短语如下:
{CJK}UTF8gbsn 短语: "<|PHRASE|>"
{CJK}UTF8gbsn 该短语向你揭示了用户或与用户相关的人或事物的,新的隐私敏感信息
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step)。
{CJK}UTF8gbsn 并总结短语泄露的具体隐私敏感信息
{CJK}UTF8gbsn 然后,输出如下格式的JSON:
“‘json
    "phrase": {CJK}UTF8gbsn 短语 ,
    "privacy information": {CJK}UTF8gbsn 隐私信息
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:"我正在准备回家,地址是上海市闵行区东川路800号。"
{CJK}UTF8gbsn 短语:"上海市闵行区东川路800号"
{CJK}UTF8gbsn JSON输出:
“‘json
    "phrase": "{CJK}UTF8gbsn 上海市闵行区东川路800号" ,
    "privacy information": "{CJK}UTF8gbsn 用户的家庭地址是上海市闵行区东川路800号。"
…… (other examples)
Table 18: Data samples with privacy phrase and privacy information. The samples are come from one raw dialogue.
Raw data examples
[…]
"id": "ShareGPT-pKNQqpRRE1",
"conversation": [
        "user": "Can you please list the operators know to work for ArcelorMittal in the world and who are not OBS?",
        "assistant": "As an AI language model, I don’t have access to up-to-date information ……",
        "user": "David Glijer is a key target for us. How would you recommand to engage him?",
        "assistant": "When engaging with David Glijer, it’s important to understand ……",
Annotated data examples of our dataset
[…]
"id": "ShareGPT-pKNQqpRRE1",
"conversation": [
        "user": "Can you please list the operators know to work for ArcelorMittal in the world and who are not OBS?",
        "assistant": "As an AI language model, I don’t have access to up-to-date information ……",
        "privacy": []
        "user": "David Glijer is a key target for us. How would you recommand to engage him?",
        "assistant": "When engaging with David Glijer, it’s important to understand ……",
        "privacy": [
                "phrase": "David Glijer",
                "privacy information": "David Glijer is identified as a key target of interest, suggesting a strategic focus in a business context."
                "phrase": "key target for us",
                "privacy information": "David Glijer is an important person of interest or focus for the user’s objectives, implying a strategic or professional engagement."
                "phrase": "engage him",
                "privacy information": "The user has plans to strategically engage or interact with David Glijer, indicating he is a subject of interest."
Table 19: Prompt template for zero-shot generation of privacy leakage classification task in English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for zero-shot generation (English).
[…]
Determine whether the query leaks the privacy information of the user, or user related people or things.
@@
“‘json
    "judgment": true or false
The user’s QUERY is
"<|QUERY|>"
Prompts for zero-shot generation (Chinese).
[…]
{CJK}UTF8gbsn 判断用户请求是否泄露了用户、用户相关人员或事物的隐私敏感信息。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。
{CJK}UTF8gbsn 以下是格式:
“‘json
    "judgment": true or false
{CJK}UTF8gbsn 用户的请求是:
"<|QUERY|>"
Table 20: Prompt template for zero-shot generation of privacy phrase extraction task in English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for zero-shot generation (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
@@
“‘json
["phrase1", "phrase2", …]
The user’s QUERY is
"<|QUERY|>"
Prompts for zero-shot generation (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性
{CJK}UTF8gbsn 隐私短语应当有具体的指代,而不是一个泛指的词汇。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是格式:
“‘json
{CJK}UTF8gbsn ["短语1", "短语2", …]
{CJK}UTF8gbsn 用户的请求是:
"<|QUERY|>"
Table 21: Prompt template for zero-shot generation of privacy information summarization task in English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for zero-shot generation (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
@@
“‘json
        "phrase": PHRASE-1,
        "privacy information": PRIVACY-INFORMATION-1
The user’s QUERY is
"<|QUERY|>"
Prompts for zero-shot generation (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
{CJK}UTF8gbsn 然后,对于每个提取的短语,总结该短语泄露的具体隐私信息
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是格式:
“‘json
{CJK}UTF8gbsn         "phrase": "短语",
{CJK}UTF8gbsn         "privacy information": "具体的隐私信息"
{CJK}UTF8gbsn 用户的请求是:
"<|QUERY|>"
Table 22: Prompt template for in-context learning of privacy leakage classification task in English and Chinese dataset, where <|INPUT|> is the input of user query and <|CASE|>is structured examples from training set.
Prompts for in-context learning (English).
[…]
Determine whether the query leaks the privacy information of the user, or user related people or things.
@@
"<|QUERY|>"
Prompts for in-context learning (Chinese).
[…]
{CJK}UTF8gbsn 判断用户请求是否泄露了用户、用户相关人员或事物的隐私敏感信息。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。
{CJK}UTF8gbsn 以下是一些例子:
{CJK}UTF8gbsn <|CASE|>例子结束。
{CJK}UTF8gbsn 用户当前的请求是:
"<|QUERY|>"
Table 23: Prompt template for in-context learning of privacy phrase extraction task in English and Chinese dataset, where <|INPUT|> is the input of user query and <|CASE|>is structured examples from training set.
Prompts for in-context learning (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.

### v2

@@ Appendix F Ethical Statement
[privacy category 1, privacy category 2, ... ]
Prompts for categories deduplication (Chinese). 给定以下¡BLOCK-SIZE¿个隐私类别及每个隐私类别对应的示例短语列表:
¡—INPUT—¿
隐私类别进行去重,如果隐私类别A和隐私类别B具有相同的含义,只保留类别A
让我们一步一步地思考,最后输出一个json列表。
以列表格式输出去重后的新的隐私类别,例如:
[ 隐私类别1, 隐私类别2 , ... ]
Figure 11. Prompt template for categories deduplication in step 2 for English and Chinese dataset, where ¡—INPUT—¿ is the input of privacy categories and corresponding phrase examples, and ¡—BLOCK-SIZE—¿ is the number of input categories.
Prompts for privacy phrase extraction (English). A user poses a query
”¡—INPUT—¿” to you.
@@
Give your think steps. The reasoning must depend on the direct evidence contained in the query and cannot be based on assumption.
@@
Privacy Categories: ¡—CATEGORIES—¿
@@
‘‘‘json { "phrase 1": "privacy category 3", "phrase 2": "privacy category 4" } ’’’ Here are some examples
User’s Query: ”I’m a teacher in China, majoring in computer science. I will go back home to London.”
Privacy Categories: [[”Name”, ”Age”, ”Gender”, ”Job”, ”Subject”, ”Location”]
@@
‘‘‘json { "teacher": "Job", "computer science": "Subject", "China": "Location", "London": "Location" } ’’’ …… (other examples)
Prompts for privacy phrase extraction (Chinese). 用户向你提出一个请求
”¡—INPUT—¿”
从请求的内容中提取短语,该短语明确揭示用户的,或与用户相关的人或事物的隐私敏感信息。
短语必须与请求内容中出现的完全一致。不要更改短语的格式或措辞,即拼写错误、空格数量、大小写等,都必须与请求内容完全相同。
给出你的思考步骤(Think Step by Step)。推理必须依赖于请求中包含的直接证据,不能基于假设
然后,将每个提取的短语与以下预定义列表中最合适的隐私类别匹配,并且如果提取的短语未与任何给定的隐私类别匹配,则应将其排除。
隐私类别:¡—CATEGORIES—¿
结果输出应为json格式,格式如下:
‘‘‘json { "短语1": "隐私类别3", "短语2": "隐私类别4" } ’’’ 以下是一些示例
用户的请求:”我是中国的一名教师,主修计算机科学。我将回家到伦敦。”
隐私类别:[”姓名”, ”年龄”, ”性别”, ”职业”, ”专业”, ”地点”]
JSON输出:
‘‘‘json { "教师": "职业", "计算机科学": "专业", "中国": "地点", "伦敦": "地点" } ’’’ …… (other examples)
Figure 12. Prompt template for privacy phrase extraction with categories in step 3 for English and Chinese datasets, where ¡—INPUT—¿ is the input of the user query and ¡—CATEGORIES—¿ is the input of privacy categories.
Prompts for privacy phrase deduplication (English).. Given a user’s query and privacy phrases extracted from the query,
Query: ”¡—INPUT—¿”
Privacy Phrases: ¡—PHRASES—¿
@@
[Rule 2] If privacy phrase A is part of privacy phrase B, keep phrase which has more information and is clearer.
@@
‘‘‘json [new privacy phrase 1, new privacy phrase 2]\\ ’’’ Here is an example:
Query: ”We are happy and plan to play football.”
Privacy Phrases: [”happy”, ”We are happy”, ”plan to play football”, ”football”, ”play football”]
@@
‘‘‘json ["happy", "play football"]\\ ’’’
Prompts for privacy phrase deduplication (Chinese). 给定一个用户请求(Query)和从请求中提取的隐私短语(Phrase),
用户请求: ”¡—QUERY—¿ ”
隐私短语: ¡—PHRASES—¿
一步一步地思考(Think Step by Step),严格按照给定的规则对隐私短语进行去重
[规则1] 如果隐私短语A和隐私短语B具有相同的含义,保留更简洁明了的短语。
[规则2] 如果隐私短语A是隐私短语B的一部分,保留信息量更大且清晰的短语。
给出你的思考步骤。然后以json列表格式输出去重后的新隐私短语,例如:
‘‘‘json [ 新隐私短语1, 新隐私短语2 ] ’’’ 以下是一个示例:
用户的请求(Query):”我们很开心,计划去踢足球。”
隐私短语:[’开心’, ’我们很开心’, ’计划去踢足球’, ’足球’, ’踢足球’]
JSON输出:
‘‘‘json [ "开心", "踢足球" ] ’’’
Figure 13. Prompt template for privacy phrase deduplication in step 3 for English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—PHRASES—¿ is the input of the extracted privacy phrase.
Prompts for filtering privacy phrase with rule 1 (English). A user poses a query
”¡—INPUT—¿”
@@
Give your think steps. The reasoning must depend on the direct evidence contained in the query and cannot be based on assumption.
@@
‘‘‘json { "reason": reason for your judgment, "judgment": true or false } ’’’ Here are some examples:
User’s Query: ”We plan to play football with Ross.”
Phrase: ”play football with Ross”
@@
‘‘‘json { "reason": "Playing football is the users plan that is related to the user. Ross is also related to the user.", "judgment": true } ’’’ …… (other examples)
Prompts for filtering privacy phrase with rule 1 (Chinese). 用户向你提出一个请求
”¡—INPUT—¿”
并给定用户的真实的身份标识
给定从请求中提取出短语如下:
短语:” ¡—PHRASE—¿ ”
一步一步地思考(Think Step by Step),判断该短语是否满足以下规则
[规则] 该短语在用户请求的上下文中直接与用户或与用户相关的人或事物相关联。
给出你的思考步骤。推理必须依赖于请求中包含的直接证据,不能基于假设
然后,输出如下格式的json:
‘‘‘json { "reason" : 你判断的理由, "judgment" : true 或 false } ’’’ 以下是一些示例:
用户的请求:”我们计划和罗斯一起踢足球。”
短语:”和罗斯一起踢足球”
JSON输出:
‘‘‘json { "reason" : "踢足球是用户的计划,与用户相关。罗斯也与用户相关。", "judgment" : true } ’’’ …… (other examples)
Figure 14. Prompt template for filtering privacy phrase with rule 1 in step 3 for English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—PHRASE—¿ is the input of the phrase.
Prompts for filtering privacy phrase with rule 2 (English). A user poses a query
”¡—INPUT—¿”
@@
[Rule] the phrase should not be a general phrase that provides NO information related to the user based on the context of the user’s query.
Give your think steps. The reasoning must depend on the direct evidence contained in the query and cannot be based on assumption.
@@
‘‘‘json { "reason": reason for your judgment, "judgment": true or false } ’’’ Here are some examples:
User’s Query: ”I want to go to a place.”
Phrases: ”place”
@@
‘‘‘json { "reason": "The place term has no clear reference.", "judgment": false } ’’’ …… (other examples)
Prompts for filtering privacy phrase with rule 2 (Chinese). 用户向你提出一个请求
”¡—INPUT—¿”
并给定用户的真实的身份标识
给定从请求中提取出短语如下:
短语:” ¡—PHRASE—¿ ”
一步一步地思考(Think Step by Step),判断该短语是否满足以下规则
[规则] 该短语不应是一个在用户请求上下文中不提供与用户相关信息的一般短语。
给出你的思考步骤。推理必须依赖于查询中包含的直接证据,不能基于假设
然后,输出如下格式的json:
‘‘‘json { "reason": 你判断的理由 , "judgment": true or false } ’’’ 以下是一些示例:
用户的请求:”我想去一个地方。”
短语:”地方”
JSON输出:
‘‘‘json { "reason": "’地方’一词没有明确的指代。" , "judgment": false } ’’’ …… (other examples)
Figure 15. Prompt template for filtering privacy phrase with rule 2 in step 3 for English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—PHRASE—¿ is the input of the phrase.
Prompts for privacy information annotation (English). A user poses a query
”¡—INPUT—¿”
@@
‘‘‘json { "phrase": PHRASE, "privacy information": PRIVACY-INFORMATION } ’’’ Here are some examples:
User’s Query: ”I am preparing to go back home to Minhang District, Shanghai.”
Phrase: ”Minhang District, Shanghai”
@@
‘‘‘json { "phrase": "No. 800, Dongchuan Road, Minhang District, Shanghai", "privacy information": "The users home address is No. 800, Dongchuan Road, Minhang District, Shanghai." } ’’’ …… (other examples)
Prompts for privacy information annotation (Chinese). 用户向你提出一个请求
”¡—INPUT—¿”
并给定用户的真实的身份标识
给定从请求中提取出短语如下:
短语: ”¡—PHRASE—¿”
该短语向你揭示了用户或与用户相关的人或事物的,新的隐私敏感信息
一步一步地思考(Think Step by Step)。
总结短语泄露的具体隐私敏感信息
然后,输出如下格式的JSON:
‘‘‘json { "phrase": 短语 , "privacy information": 隐私信息 } ’’’ 以下是一些示例:
用户的请求:”我正在准备回家,地址是上海市闵行区东川路800号。”
短语:”上海市闵行区东川路800号”
JSON输出:
‘‘‘json { "phrase": "上海市闵行区东川路800号" , "privacy information": "用户的家庭地址是上海市闵行区东川路800号。" } ’’’ …… (other examples)
Figure 16. Prompt template for privacy information annotation in step 4 for English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—PHRASE—¿ is the input of the phrase.
Raw data examples. "id": "ShareGPT-pKNQqpRRE1", "conversation": [ { "user": "Can you please list the operators know to work for ArcelorMittal in the world and who are not OBS?", "assistant": "As an AI language model, I dont have access to up-to-date information ......", }, { "user": "David Glijer is a key target for us. How would you recommand to engage him?", "assistant": "When engaging with David Glijer, its important to understand ......", } ]
Annotated data examples of our dataset. "id": "ShareGPT-pKNQqpRRE1", "conversation":[ { "user": "Can you please list the operators know to work for ArcelorMittal in the world and who are not OBS?", "assistant": "As an AI language model, I dont have access to up-to-date information ......", "privacy": [] }, { "user": "David Glijer is a key target for us. How would you recommand to engage him?", "assistant": "When engaging with David Glijer, its important to understand ......", "privacy": [ { "phrase": "David Glijer", "privacy information": "David Glijer is identified as a key target of interest, suggesting a strategic focus in a business context." }, { "phrase": "key target for us", "privacy information": "David Glijer is an important person of interest or focus for the users objectives, implying a strategic or professional engagement." }, { "phrase": "engage him", "privacy information": "The user has plans to strategically engage or interact with David Glijer, indicating he is a subject of interest." } ] } ]
Figure 17. Data samples with privacy phrase and privacy information. The samples come from one raw dialogue.
Prompts for zero-shot generation (English). Determine whether the query leaks the privacy information of the user, or user related people or things.
@@
‘‘‘json { "judgment": true or false } ’’’ The user’s QUERY is
”¡—QUERY—¿”
Prompts for zero-shot generation (Chinese). 判断用户请求是否泄露了用户、用户相关人员或事物的隐私敏感信息。
隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
请一步一步思考(Think step-by-step)。
给出理由并在最后,以JSON格式输出答案。
以下是格式:
‘‘‘json { "judgment": true or false } ’’’ 用户的请求是:
”¡—QUERY—¿”
Figure 18. Prompt template for zero-shot generation of privacy leakage classification task in English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query.
Prompts for zero-shot generation (English). Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
@@
‘‘‘json ["phrase1", "phrase2", ...] ’’’ The user’s QUERY is
”¡—QUERY—¿”
Prompts for zero-shot generation (Chinese). 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性
隐私短语应当有具体的指代,而不是一个泛指的词汇。
请一步一步思考(Think step-by-step)。
给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
以下是格式:
‘‘‘json [ "短语1", "短语2", ... ] ’’’ 用户的请求是:
”¡—QUERY—¿”
Figure 19. Prompt template for zero-shot generation of privacy phrase extraction task in English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query.
Prompts for zero-shot generation (English). Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
@@
‘‘‘json [ { "phrase": PHRASE-1, "privacy information": PRIVACY-INFORMATION-1 } ] ’’’ The user’s QUERY is
”¡—QUERY—¿”
Prompts for zero-shot generation (Chinese). 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
然后,对于每个提取的短语,总结该短语泄露的具体隐私信息
请一步一步思考(Think step-by-step)。
给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
以下是格式:
‘‘‘json [ { "phrase": "短语", "privacy information": "具体的隐私信息" } ] ’’’ 用户的请求是:
”¡—QUERY—¿”
Figure 20. Prompt template for zero-shot generation of privacy information summarization task in English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query.
Prompts for in-context learning (English). Determine whether the query leaks the privacy information of the user, or user related people or things.
@@
”¡—QUERY—¿”
Prompts for in-context learning (Chinese). 判断用户请求是否泄露了用户、用户相关人员或事物的隐私敏感信息。
隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
请一步一步思考(Think step-by-step)。
给出理由并在最后,以JSON格式输出答案。
以下是一些例子:
¡—CASE—¿例子结束。
用户当前的请求是:
”¡—QUERY—¿”
Figure 21. Prompt template for in-context learning of privacy leakage classification task in English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—CASE—¿is structured examples from the training set.
Prompts for in-context learning (English). Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.